OpenStudy (anonymous):

Which phrase matches the algebraic expression below? 2(x − 7) + 10 Two times the sum of x and seven plus ten Two times the difference of x and seven plus ten Two times x minus seven plus ten Two times x minus the sum of seven and ten

OpenStudy (anonymous):

Two times the difference of x and seven plus ten

OpenStudy (anonymous):

The difference of x and seven = x - 7 Two times the difference of x and seven = 2(x - 7) Two times the difference of x and seven plus ten = 2(x - 7) + 10